OpenStudy (anonymous):

Find the product of (x − 3)2. x2 + 6x + 9 x2 − 9 x2 − 6x + 9 x2 + 9

OpenStudy (anonymous):

Same thing?

OpenStudy (anonymous):

x^2 - 9?

Parth (parthkohli):

\(\Large \color{Midnightblue}{\rightarrow (a - b)^2 = a^2 + b^2 - 2ab }\)

Parth (parthkohli):

Basically, \(\Large \color{Midnightblue}{\rightarrow (a \pm b)^2 = a^2 + b^2 \pm 2ab }\)

OpenStudy (anonymous):

x^2 - 9 - 2x^3?

OpenStudy (anonymous):

x^2 + 9 - 6?

Parth (parthkohli):

- 6x*

Parth (parthkohli):

\(x^2 + 9 - 6x\)

OpenStudy (anonymous):

Explain please...

OpenStudy (anonymous):

Why - 6x?

Parth (parthkohli):

\(\Large \color{Midnightblue}{\rightarrow (x - 3)^2 = x^2 + 3^2 - 2(x * 3) }\) \(\Large \color{Midnightblue}{\rightarrow (x - 3)^2 = x^2 + 3^2 - 6x }\)

OpenStudy (anonymous):

I see.

OpenStudy (anonymous):

But that's not one of my options..

OpenStudy (anonymous):

@ParthKohli That's not one of my options..

OpenStudy (anonymous):

@ParthKohli That's not one of my options..

Parth (parthkohli):

Just simplify it.

OpenStudy (anonymous):

Umm... How?

Parth (parthkohli):

\(\Large \color{Midnightblue}{\rightarrow x^2 + 9 -6x }\)

Parth (parthkohli):

\(3^2 = 9\)

OpenStudy (anonymous):

None of them are my options either..

Parth (parthkohli):

It's the third one!

OpenStudy (anonymous):

Try this using distributive property . (x-3)(x-3) x*x+(-3)(-3)+(-3)x+(-3x). x^2+9-6x

Parth (parthkohli):

My answer was right, but it wasn't according to the degrees.